# From Tools to Teammates: Evaluating LLMs in Multi-Session Coding Interactions

### v1

@@ Abstract
Marco Del Tredici started this project while working at Cohere. Lucas Weber is now at Fraunhofer IIS and Alberto Testoni at Amsterdam UMC.
Large Language Models (LLMs) are increasingly used in working environments for a wide range of tasks, excelling at solving individual problems in isolation. However, are they also able to effectively collaborate over long-term interactions? To investigate this, we introduce MemoryCode, a synthetic multi-session dataset designed to test LLMs’ ability to track and execute simple coding instructions amid irrelevant information, simulating a realistic setting. While all the models we tested handle isolated instructions well, even the performance of state-of-the-art models like GPT-4o deteriorates when instructions are spread across sessions. Our analysis suggests this is due to their failure to retrieve and integrate information over long instruction chains. Our results highlight a fundamental limitation of current LLMs, restricting their ability to collaborate effectively in long interactionsThe code and data are available at https://github.com/for-ai/MemoryCode..
@@
Nathanaël Carraz Rakotonirina1, Mohammed Hamdy5, Jon Ander Campos3, Lucas Weber1, Alberto Testoni2, Marzieh Fadaee4, Sandro Pezzelle2, Marco Del Tredici 1Universitat Pompeu Fabra, 2University of Amsterdam, 3Cohere, 4Cohere For AI, 5Cohere For AI Community
@@ 2.2 Long-term Dialogue Evaluation
Most similar to our work, MMMT-IF (Epstein et al., 2024) extends multi-turn and multi-modal datasets to measure instruction-following abilities and shows that the main challenge for LLMs is not in following instructions, but rather in retrieving instructions. In multi-session dialogues, many datasets were also created synthetically; for example, Conversation Chronicles (Jang et al., 2023), which includes 200K conversations of about 5 sessions each, or LoCoMo (Maharana et al., 2024), a multi-modal dataset based on a framework that leverages personas and temporal event graphs. Kim et al. (2024a) and Kim et al. (2024b) utilized movie scripts to construct complex multi-session dialogue datasets having, for example, multi-party conversations and shared memories between speakers. Most of the works mentioned above focus on expanding the number of turns and sessions or on introducing more complex tasks that are challenging for LLMs. Similarly, our work evaluates the performance of LLMs in multi-turn/session dialogues, but with very simple tasks and more focus on practical, real-world settings in which information is constantly changing.
@@ 4.1 Models
We test several recent LLMs on our benchmark, namely, three versions of Llama-3.1 (8B-Instruct, 70B-Instruct, and 405B-Instruct; Dubey et al., 2024), Command R+ (Cohere, 2024b), and GPT-4o (OpenAI, 2024). Our model selection includes both proprietary and open-weights models, covering a broad range of model sizes. This provides us with a comprehensive overview of how various types of LLMs perform on our dataset. We note that all the models have been trained on code and tested on Python coding benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). The details to reproduce the results are provided in Appendix C.
@@ 5 Results
Figure 3: Average Instruction, Session, and History scores per model. For the latter, ‘short’ includes dialogue histories with less than 15 sessions, ‘long’ those with 16 to 100 sessions.
@@
As shown in Figure 3, all models achieve high or very high accuracy on this setup. This aligns with our goals of having relatively easy instructions. In particular, all the large models approach or exceed 0.9 accuracy—reported in the $[0,1]$ scale. While the results of Llama-8B are lower, they still show that even a small, non-specialized model achieves good performance, confirming the easiness of the task.
As a sanity check, we run a set of experiments in which we do not prompt the models with the necessary instructions (e.g., use CamelCase), but directly run the evaluation (in this case, we check if CamelCase was used).555 Due to budget limits, we only used Command R+ for this experiment and for the analysis in Section 6.2. We expect the results to be representative of all other models’ behavior. This setup verifies that models do not solve MemoryCode through their default behavior. Models fail spectacularly, achieving an extremely low average accuracy (consistently lower than 0.01), confirming that the instructions we provide are crucial to executing the tasks correctly.
@@
The performance in this setup is very similar to Instruction for the larger Llama models and GPT-4o, indicating that these models have no difficulties at retrieving the relevant information in a single session. Command R+ shows a larger drop of 0.22 (25% relative drop compared to Instruction), while Llama-3.1-8B shows a major drop of 0.34 (48%), which indicates its inability to retrieve relevant information across multiple turns.
@@
Things change dramatically in this setup, with a degradation in performance across the board already for ‘short’ dialogue histories. In particular, GPT-4o shows a drop of 0.14, a value that increases for the other models: 0.47 (67% relative drop compared to Instruction) for Llama-3.1-8B and 0.48 (54%) for Command R+. These results indicate that as the number of sessions increases, even the best-in-class models have difficulties in identifying and applying the relevant instructions.
A more dramatic drop is observed in the ‘long’ setup. Here, all the models struggle, with the best GPT-4o only achieving 0.30 accuracy, which indicates a relative drops of 61% from the ‘short’ setup and of 67% from Instruction. The drop is even more significant for the other models: the performance of Llama-3.1-405B drops by 78% compared to Instruction, Command R+ by 87%. Crucially, this happens even though the tasks on which models are evaluated are identical to those in the Instruction and Session setups, where the models achieved nearly-perfect accuracy. The difference in performance, hence, is to be ascribed to models’ inability to retrieve and reason over relevant pieces of information present in their input.
@@ 6.1 Effect of Number of Sessions
This weakness could be due to limitations in retrieving the relevant information from the dialogue history, reasoning about a chain of instructions and updates, or both. Below, we shed light on this issue.
6.2 Retrieval or Reasoning Problem?
If the poor performance in the History setup was due to retrieval, then passing models only the full chain of instructions and updates—without any intervening irrelevant text—should solve the issue. Vice versa, if the issue was about reasoning over such a chain, they should still perform poorly. We test these assumptions by feeding Command R+ with only the entire chain of instructions needed to solve a task. We name this setup Instructions-chain. As shown in Figure 5, the trend is strikingly similar to the one observed in History, with the model still struggling even if only the relevant information is provided, with no dialogue history. This indicates that models’ drop in performance is mainly due to their inability to reason compositionally over a sequence of instructions.
@@ 6.3 Effect of Instruction Updates
After uncovering models’ reasoning limitations, we now explore whether they are due to the inability to cope with instruction updates. We hypothesize that the higher the number of updates an instruction goes through, the harder it is for the model to keep track of such updates and eventually adhere to the updated instruction when performing a task.
We define the update rank of an instruction as the number of times the instruction is updated throughout the dialogue history, for both ‘short’ and ‘long’ setups. An update rank of 0 means that the instruction was never updated. Figure 6 reports for each update rank the average score, computed across all models, of the instructions with that rank. We observe that the score of an instruction decreases as its number of updates increases. This result is consistent with our hypothesis that models struggle to incrementally update their instruction representations through multiple reasoning steps.
@@ 6.4 Instruction Difficulty
We finally assess if, besides the updates, the very nature of each instruction makes it more challenging than others. For this, we compute the average per-instruction score of the best-performing model, GPT-4o, for both the initial instruction and their updates. The results reveal notable variations across instructions: In particular, instructions and updates that are less common in practice—like including a digit in object names—tend to obtain lower scores compared to more common ones such as using docstrings or annotations. Figure 8 and 9 in Appendix E report the results of this analysis. Overall, this suggests that the models’ ability to retrieve and update relevant instructions is modulated by their characteristics and that familiarity is likely to play a role. An example (from GPT-4o) is shown in Table 3. In this case, the model correctly follows the instruction on comments, but not those about attribute and variable names.
@@ 7 Conclusions
Overall, our results show a severe limitation of current LLMs. The inability to keep track of simple information as the interaction with a human unfolds effectively hinders their adoption in real-world scenarios and restricts their usage to addressing single, self-contained problems. We argue that effective long-term collaboration cannot be achieved by further scaling model and input context window sizes. Rather, we believe that new mechanisms to handle and retrieve from long-term memory need to be developed. MemoryCode contributes to this challenging and yet crucial goal, by providing a robust benchmark for developing and testing such methods.
@@ Acknowledgments
UPF was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101019291). During his affiliation with UvA, Alberto Testoni was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 819455, PI R. Fernández). This paper reflects the authors’ view only, and the funding agency is not responsible for any use that may be made of the information it contains.
@@ Appendix H Ethics statement
Figure 8: Average per-instruction insertion scores for GPT-4o
Figure 9: Average per-instruction update scores for GPT-4o

### v2

@@ Abstract
Marco Del Tredici started this project while working at Cohere and Alberto Testoni while working at the University of Amsterdam. Lucas Weber is now at Fraunhofer IIS.
Large Language Models (LLMs) are increasingly used in working environments for a wide range of tasks, excelling at solving individual problems in isolation. However, are they also able to effectively collaborate over long-term interactions? To investigate this, we introduce MemoryCode, a synthetic multi-session dataset designed to test LLMs’ ability to track and execute simple coding instructions amid irrelevant information, simulating a realistic setting. While all the models we tested handle isolated instructions well, even state-of-the-art models like GPT-4o and the reasoning model DeepSeek-R1 show degraded performance when instructions are spread across sessions. Our analysis suggests this is due to their failure to retrieve and integrate information over long instruction chains. Our results highlight a fundamental limitation of current LLMs, restricting their ability to collaborate effectively in long interactionsThe code and data are available at https://github.com/Cohere-Labs-Community/MemoryCode..
@@
Nathanaël Carraz Rakotonirina1, Mohammed Hamdy5, Jon Ander Campos3, Lucas Weber1, Alberto Testoni6, Marzieh Fadaee4, Sandro Pezzelle2, Marco Del Tredici 1Universitat Pompeu Fabra, 2University of Amsterdam, 3Cohere, 4Cohere Labs, 5Cohere Labs Community, 6Amsterdam UMC
@@ 2.2 Long-term Dialogue Evaluation
Most similar to our work, MMMT-IF (Epstein et al., 2024) extends multi-turn and multi-modal datasets to measure instruction-following abilities and shows that the main challenge for LLMs is not in following instructions, but rather in retrieving it. Unlike MMMT-IF, our dataset was generated entirely from scratch and is specifically designed for real-world conversational settings that feature constantly evolving instructions. In multi-session dialogues, many datasets were also created synthetically; for example, Conversation Chronicles (Jang et al., 2023), which includes 200K conversations of about 5 sessions each, or LoCoMo (Maharana et al., 2024), a multi-modal dataset based on a framework that leverages personas and temporal event graphs. Kim et al. (2024a) and Kim et al. (2024b) utilized movie scripts to construct complex multi-session dialogue datasets having, for example, multi-party conversations and shared memories between speakers. Most of the works mentioned above focus on expanding the number of turns and sessions or on introducing more complex tasks that are challenging for LLMs. Similarly, our work evaluates the performance of LLMs in multi-turn/session dialogues, but with very simple tasks and more focus on practical, real-world settings in which information is constantly changing.
@@ 4.1 Models
We test several recent LLMs on our benchmark, namely, three versions of Llama-3.1 (8B-Instruct, 70B-Instruct, and 405B-Instruct; Dubey et al., 2024), Command R+ (Cohere, 2024b), Command A (Cohere et al., 2025), GPT-4o (OpenAI, 2024), DeepSeek-V3 (Liu et al., 2024), DeepSeek-R1 (Guo et al., 2025). Our model selection includes both proprietary and open-weights models, as well as reasoning and non-reasoning models, and a broad range of model sizes. This provides us with a comprehensive overview of how various types of LLMs perform on our dataset. We note that all the models have been trained on code and tested on Python coding benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). The details to reproduce the results are provided in Appendix C.
@@ 5 Results
Figure 3: Average Instruction, Session, and History scores per model. For the latter, ‘short’ includes dialogue histories with less than 15 sessions, ‘long’ those with 16 to 100 sessions. Results include 95% confidence intervals.
@@
As shown in Figure 3, all models achieve high or very high accuracy on this setup. This aligns with our goals of having relatively easy instructions. In particular, all the large models approach or exceed 0.9 accuracy score—reported in the $[0,1]$ scale— with DeepSeek-V3 achieving a perfect. While the results of Llama-8B are lower, they still show that even a small, non-specialized model achieves good performance, confirming the easiness of the task.
As a sanity check, we run a set of experiments in which we do not prompt the models with the necessary instructions (e.g., use CamelCase), but directly run the evaluation (in this case, we check if CamelCase was used).555 Due to budget limits, we only used Command R+ for this experiment and for the analysis in Section LABEL:subsec:retrieval_or_reasoning_problem. We expect the results to be representative of all other models’ behavior. This setup verifies that models do not solve MemoryCode through their default behavior. Models fail spectacularly, achieving an extremely low average accuracy (consistently lower than 0.01), confirming that the instructions we provide are crucial to executing the tasks correctly.
@@
The performance in this setup is very similar to Instruction for the larger Llama models, DeepSeek models, Command A, and GPT-4o, indicating that these models have no difficulties at retrieving the relevant information in a single session. Command R+ shows a larger drop of 0.22 (25% relative drop compared to Instruction), while Llama-3.1-8B shows a major drop of 0.34 (48%), which indicates its inability to retrieve relevant information across multiple turns.
@@
Things change dramatically in this setup, with a degradation in performance across the board already for ‘short’ dialogue histories. In particular, GPT-4o shows a drop of 0.14. A similar decrease is observed in the larger LLaMA models, DeepSeek models, and Command R. The drop in score is even greater for the other models: 0.47 (67% relative drop compared to Instruction) for Llama-3.1-8B and 0.48 (54%) for Command R+. These results indicate that as the number of sessions increases, even the best-in-class models have difficulties in identifying and applying the relevant instructions.
A more dramatic drop is observed in the ‘long’ setup. Here, all the models struggle, with GPT-4o, the best non-reasoning model, only achieving 0.30 accuracy, which indicates a relative drops of 61% from the ‘short’ setup and of 67% from Instruction. While DeepSeek-R1 significantly improves over the performance of GPT-4o, it still shows a drop of 0.56 points compared to the performance on the Instruction setup. These results confirm that the main challenge of the benchmark is not merely following individual instructions but reasoning over a sequence of instructions. Moreover, they highlight that even newer models equipped with reasoning abilities struggle on this benchmark, suggesting that complementary or alternative capabilities are needed to solve it effectively. The drop is even more significant for the other models: the performance of Llama-3.1-405B drops by 78% compared to Instruction, Command R+ by 87%. Crucially, this happens even though the tasks on which models are evaluated are identical to those in the Instruction and Session setups, where the models achieved nearly-perfect accuracy. The difference in performance, hence, is to be ascribed to models’ inability to retrieve and reason over relevant pieces of information present in their input.
@@ 6.1 Effect of Number of Sessions
This weakness could be due to limitations in retrieving the relevant information from the dialogue history, compositionally applying instructions from the chain, inferring their latest updates, or any combination of these factors. Below, we shed light on this issue.
6.2 Retrieval or Multi-Instruction-following Problem?
If the poor performance in the History setup was due to retrieval, then passing models only the full chain of instructions and updates—without any intervening irrelevant text—should solve the issue. Vice versa, if the issue stemmed from difficulties in understanding and jointly applying multiple instructions, they should still perform poorly. We test these assumptions by feeding Command R+ with only the entire chain of instructions needed to solve a task. We name this setup Instructions-chain. As shown in Figure 5, the trend is strikingly similar to the one observed in History, with the model still struggling even if only the relevant information is provided, with no dialogue history. This indicates that models’ drop in performance is mainly due to their inability to reason compositionally over a sequence of instructions.
@@ 6.3 Effect of Instruction Updates
After delineating the contribution of retrieval and multiple-instruction-following to the performance degradation, we now examine the role of instruction updates.
We define the update rank of an instruction as the number of times the instruction is updated throughout the dialogue history, for both ‘short’ and ‘long’ setups. An update rank of 0 means that the instruction was never updated. The maximum number of updates in our dataset is 7.
The number of updates increases with number of sessions: to control for the latter, we compute the scores for the different update ranks for each number of session independently.
Figure 6 reports the result for each update rank obtained with GPT-4o. The results show that the score is stable across different update ranks.
This suggests that the poor performance on this dataset is primarily due to the model’s difficulty in jointly following multiple rules, rather than its ability to update a single instruction.
@@ 6.4 Instruction Difficulty
We finally assess if, besides the updates, the very nature of each instruction makes it more challenging than others. For this, we compute the average per-instruction score of the best-performing model, GPT-4o, for both the initial instruction and its updates. The results reveal notable variations across instructions: In particular, instructions and updates that are less common in practice—like those including a digit in object names—tend to obtain lower scores compared to more common ones, such as using docstrings or annotations. Figure 8 and 9 in Appendix E report the results of this analysis. Overall, this suggests that the models’ ability to retrieve and update relevant instructions is modulated by their characteristics and that familiarity is likely to play a role. An example (from GPT-4o) is shown in Table 3. In this case, the model correctly follows the instruction on comments, but not those about attribute and variable names.
We note that instruction difficulty naturally interacts with another key aspect of long-term interactions: the degree of compositionality among the pieces of information provided by the mentor. Besides overwriting prior instructions or appending complementary ones, mentors may issue instructions that combine in compositional ways. For instance, one instruction might require that any line of code using parallel computing libraries (e.g., NumPy or Torch) be followed by a comment, while another might call for merging comments from adjacent lines if the associated code is highly correlated. Although we do not explicitly analyze such cases here, we argue that our approach can be readily adapted to systematically study these scenarios in a controlled setting.
@@ 7 Conclusions
Overall, our results show a severe limitation of current LLMs. The inability to keep track of simple information as the interaction with a human unfolds effectively hinders their adoption in real-world scenarios and restricts their usage to addressing single, self-contained problems. We argue that effective long-term collaboration cannot be achieved by further scaling alone, be it model or dataset size, context window, or test-time compute. Rather, we believe that new mechanisms to handle and retrieve from long-term memory need to be developed. MemoryCode contributes to this challenging and yet crucial goal, by providing a robust benchmark for developing and testing such methods.
@@ Acknowledgments
UPF was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101019291). During his work at the University of Amsterdam, Alberto Testoni was supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 819455, PI R. Fernández). Alberto Testoni is currently funded by the project CaRe-NLP (PI Iacer Calixto) with file number NGF.1607.22.014 of the research programme AiNed Fellowship Grants, which is (partly) financed by the Dutch Research Council (NWO).This paper reflects the authors’ view only, and the funding agency is not responsible for any use that may be made of the information it contains.
@@ Appendix H Ethics statement
Figure 8: Average per-instruction insertion scores for GPT-4o. Results include 95% confidence intervals.
Figure 9: Average per-instruction update scores for GPT-4o. Results include 95% confidence intervals.